# Improving LLM Predictions via Inter-Layer Structural Encoders

### 24 Mar 2026 v1

@@ 5 Discussion and Future Work
As future work, several promising directions emerge. While we evaluated models up to 8B parameters, future work could investigate whether ILSE holds for larger architectures (e.g., Llama-3-70B) or in generative contexts, such as mitigating hallucinations through enhanced multi-layer representation mixing. Additionally, we can explore alternative mappings from layers to nodes in cayley topology, such as random or learned permutations, or task-specific mappings for particular downstream tasks.
Another interesting direction is implementing Cayley aggregation over tokens for a single layer as a token pooling method versus standard mean or last token pooling. This could be combined with layer-wise aggregation for geometric fusion at both levels. Finally, ILSE currently utilizes the structures as a task-specific fusion layer for frozen models. A promising direction is the integration of ILSE directly into the transformer backbone during pre-training. Such a ”geometry-aware” architecture could potentially yield more robust base models with improved global information flow.

### 28 Apr 2026 v2

@@ 5 Discussion and Future Work
As future work, several promising directions emerge. While we evaluated models up to 8B parameters, future work could investigate whether ILSE holds for larger architectures (e.g., Llama-3-70B) or in generative contexts, such as mitigating hallucinations through enhanced multi-layer representation mixing. Additionally, we can explore alternative mappings from layers to nodes in cayley topology, such as learned permutations, or task-specific mappings for particular downstream tasks.
Finally, ILSE currently utilizes the structures as a task-specific fusion layer for frozen models. A promising direction is the integration of ILSE directly into the transformer backbone during pre-training. Such a ”geometry-aware” architecture could potentially yield more robust base models with improved global information flow.